### v2

@@ Abstract
Randomly moving active particles can be herded into directed motion by asymmetric geometric structures. Although such a rectification process has been extensively studied due to its fundamental and technological relevance, a comprehensive understanding of active matter rectification based on single particle dynamics remains elusive. Here, by combining experiments, simulations, and theory, we study the directed transport and energetics of swimming bacteria navigating through funnel-shaped obstacles—a paradigmatic model of rectification of living active matter. We develop a microscopic parameter-free model for bacterial rectification, which quantitatively explains experimental and numerical observations and predicts the optimal geometry for the maximum rectification efficiency. Furthermore, we quantify the degree of time irreversibility and measure the extractable work associated with bacterial rectification. Our study provides quantitative solutions to long-standing questions on bacterial rectification and establishes a generic relationship between time irreversibility, particle fluxes, and extractable work, shedding light on the energetics of non-equilibrium rectification processes in living systems.
@@
This manuscript was compiled on June 11, 2024 www.pnas.org/cgi/doi/10.1073/pnas.XXXXXXXXXX
6]3
@@
Composed of independent self-propelled units, active matter provides a paradigm for the study of symmetry-breaking mechanisms and non-equilibrium dynamics in rectified living systems under well controlled conditions. The self-propulsion of active particles intrinsically breaks time-reversal symmetry, whereas the breaking of spatial inversion symmetry can be achieved independently by using asymmetric geometric structures (7, 8, 9, 10, 11, 12, 13). Notably, in seminal experiments on swimming bacteria, Austin, Chaikin and co-workers found that microfluidic devices with an array of fixed funnel-shaped obstacles spontaneously induce a concentration difference of active particles between the two sides of the array from an initially uniform active bath (14). The simplicity of the bacterial rectification process, together with the great potential in biotechnological applications such as cell sorting and trapping (14, 15, 16, 17), micro-patterning (14), spontaneous pumping (14) and cargo transport (18), has stimulated extensive studies on the rectification of active matter in recent years (19), both numerically (20, 21, 22, 23) and in experiments (14, 24, 25, 23, 26, 27, 28, 17, 24). Although it is generally acknowledged that bacterial rectification arises from bacterium-obstacle interactions (14), a quantitative understanding of the directed transport based on the microscopic dynamics of bacteria has not yet been achieved. Despite intensive research over the last two decades, important questions such as the identification of the geometry with the maximum rectification efficiency remain unresolved.
@@
Here, by combining experiments, simulations and theory, we provide a quantitative microscopic understanding of the transport and energetics of bacterial rectification—a paradigmatic model of the rectification of living active matter (19, 29). In particular, we identify three essential factors controlling bacterial rectification: (i) a universal distribution of the orientation of bacteria that swim across a linear segment such as the opening of obstacles; (ii) the realignment of bacteria along solid boundaries; (iii) bacterial wobbling. A model incorporating these factors quantitatively describes bacterial transport spanning a wide range of obstacle shapes and predicts the optimal geometry that maximizes the rectification efficiency. Furthermore, we characterize the time irreversibility of bacterial rectification in terms of its local entropy production rate (EPR) and analyze the energetics of the process. Finally, we devise a protocol to measure extractable work through a weak coupling to bacterial flux, which allows us to elucidate the intrinsic relation between three fundamental thermodynamics quantities—particle fluxes, EPR, and extractable work, the first of this kind in a single non-equilibrium process. Taken together, our study provides not only important insights into rectification processes in living systems but also a useful guideline for controlling active particles in biotechnological applications (14, 15, 16, 17, 18).
@@
Away from funnel walls, freely swimming E. coli exhibit the classic run-and-tumble motion (30). The average run length of bacteria in our quasi-2D chamber is significantly longer than the length of the funnel walls. Notably, the trajectory of a bacterium in the run phase is a 3D helix, which manifests as wobbling of the bacterial body under the 2D projection of optical microscopy (31, 32) (Movies 1 and 2). Independent of the angle of incidence, a bacterium always re-aligns and moves parallel to the funnel walls after hitting the walls (Figs. 1A and B) (14). In our minimal simulations, we model bacteria as non-interacting point particles performing run-and-tumble motions without wobbling, and incorporate the surface alignment trait of E. coli via event-driven dynamics (Materials and Methods).
@@
The agreement between our minimal simulations and experiments for $\theta<\theta_{a}$ has two important implications. First, even though bacteria-wall interactions are governed by complex steric and near-field hydrodynamic forces (33, 34, 35, 36), from the perspective of rectification, such complex interactions can be effectively condensed into one simple rule: E. coli reorient themselves parallel to the wall after a collision. Such a simple rule has been ignored in most numerical studies of bacterial rectification (20, 23), with notable exceptions (21, 22). Second, the rectification efficiency can be defined as the net flux per particle, $\eta_{r}\equiv({N}_{+}-{N}_{-})/({N}_{+}+{N}_{-})$. The non-monotonic trend in ${N}_{+}$, in combination with the nearly constant ${N}_{-}$, results in a maximum $\eta_{r}$ at $\theta=\theta_{\max}$ (Fig. 2C).
@@
The hallmark of non-equilibrium processes—with rectification in living systems as a particular example—is the presence of non-zero fluxes and, more fundamentally, time-reversal symmetry breaking (TRSB), measured by the entropy production rate (EPR), $\dot{S}$ (38, 39, 40). EPR gives a direct measure of the minimum energetic cost required to maintain a system out of equilibrium and is crucial in understanding the energetics of non-equilibrium processes.
For a system in non-equilibrium steady state (NESS), the Kullback-Leibler divergence (KLD), $\sigma$, between the probability of observing a time-forward trajectory (X) and its time-reversed counterpart ($\textbf{X}^{\textbf{R}}$) bounds $\dot{S}$ as (41, 42, 43, 44, 45, 46, 47, 48, 23, 42),
@@ Discussion
Our study explores the rectification of bacteria in the dilute limit, where the motions of bacteria are uncorrelated. How do the concepts described here extend to dense systems with correlated bacterial motions? A particularly interesting class of such systems are those displaying active turbulence (53, 54). While it has recently been shown that asymmetric obstacles placed in active turbulent baths can move persistently as autonomous engines (12, 9) and that the statistics of Lagrangian tracers in active turbulent baths carry information about time-irreversibility (55), a unifying thermodynamic framework for active turbulence is still missing, which is an interesting topic for future research.
Is the relation between local fluxes, time irreversibility, and extractable work uncovered in our study applicable to other non-equilibrium systems? Although statistical irreversibility has been widely used to quantify TRSB in various non-equilibrium phenomena (47, 48, 23, 56, 57), simultaneous measurements of time irreversibility, fluxes, and extractable work had not been achieved in any single non-equilibrium systems heretofore. We hope our results will stimulate future research in this direction, which would be of great value for illustrating general principles of non-equilibrium thermodynamics.
@@
A wild-type E. coli strain (BW 25113) genetically modified to express proteorhodopsin (a light-driven transmembrane proton pump) was used in our experiments (58). The bacterial stock was first inoculated in Terrific Broth (TB) supplemented with ampicillin in an orbital shaker at $35\degree$C for 15 hours. This culture was further diluted (1:100) in fresh TB and grown at $30\degree$C for 6-7 hours, with the addition of 1 mM isopropyl $\beta$-d-1-thiogalactopyranoside and 10 $\mu$M methanolic all-trans-retinal in the mid-log phase to trigger the expression of proteorhodopsin. The bacterial suspension was centrifuged ($800g$ for 5 min) to collect motile cells and the supernatant was resuspended in motility buffer (0.01 M potassium phosphate, 0.067 M NaCl, $10^{-4}$ M EDTA, pH 7.0). The suspension was further washed twice, and the final concentration was measured using a biophotometer via optical density at 600 nm (OD600). Bacterial concentrations in all our experiments were fixed below the threshold concentration for collective bacterial swimming (58, 53). With the supply of oxygen, the average swimming speed of bacteria was 25 $\mu$m/s.
@@
The length, width, and gap of the funnel rectifier were fixed at $30$ $\mu$m, $5$ $\mu$m, and $5$ $\mu$m, respectively. The channel depth was kept at 5-6 $\mu$m, so that bacteria were always close to the focal plane set at the middle of the channels. A bacterial suspension of concentration $8\times 10^{7}$ cells/ml was carefully introduced in the microchannels and imaged via an inverted bright-field microscope (Nikon Ti-Eclipse) using a 40×40$\times$40 × objective lens with numerical aperture (NA) 0.65. Videos were acquired at a frame rate of 20 frames per second using an Andor Zyla sCMOS camera. Bacteria were manually counted at the funnel tip for each funnel angle.

### v3

@@ Abstract
Randomly moving active particles can be herded into directed motion by asymmetric geometric structures. Although such a rectification process has been extensively studied due to its fundamental, biological, and technological relevance, a comprehensive understanding of active matter rectification based on single particle dynamics remains elusive. Here, by combining experiments, simulations, and theory, we study the directed transport and energetics of swimming bacteria navigating through funnel-shaped obstacles—a paradigmatic model of rectification of living active matter. We develop a microscopic parameter-free model for bacterial rectification, which quantitatively explains experimental and numerical observations and predicts the optimal geometry for the maximum rectification efficiency. Furthermore, we quantify the degree of time irreversibility and measure the extractable work associated with bacterial rectification. Our study provides quantitative solutions to long-standing questions on bacterial rectification and establishes a generic relationship between time irreversibility, particle fluxes, and extractable work, shedding light on the energetics of non-equilibrium rectification processes in living systems.
@@
This manuscript was compiled on June 20, 2024 www.pnas.org/cgi/doi/10.1073/pnas.XXXXXXXXXX
3]3
@@
Composed of independent self-propelled units, active matter provides a paradigm for the study of symmetry-breaking mechanisms and non-equilibrium dynamics in rectified living systems under well controlled conditions. The self-propulsion of active particles intrinsically breaks time-reversal symmetry, whereas the breaking of spatial inversion symmetry can be achieved independently by using asymmetric geometric structures (7, 8, 9, 10, 11, 12, 13). Notably, in seminal experiments on swimming bacteria, Austin, Chaikin and co-workers found that microfluidic devices with an array of fixed funnel-shaped obstacles spontaneously induce a concentration difference of active particles between the two sides of the array from an initially uniform active bath (14). The simplicity of the bacterial rectification process, together with the great potential in biotechnological applications such as cell sorting and trapping (14, 15, 16, 17), micro-patterning (14), spontaneous pumping (14) and cargo transport (18), has stimulated extensive studies on the rectification of active matter in recent years (19), both numerically (20, 21, 22, 23) and in experiments (14, 24, 25, 23, 26, 27, 28, 17, 24). Although it is generally acknowledged that bacterial rectification arises from bacterium-obstacle interactions (14), a quantitative understanding of the directed transport based on the microscopic dynamics of bacteria has not yet been achieved. Despite intensive research over the last two decades, important questions such as the identification of the geometry with the maximum rectification efficiency remain unresolved. Answers to these questions may shed light on the evolution of carnivorous plants that rectify the motion of microbes through root hairs structurally similar to funnel-shaped obstacles (29).
@@
Here, by combining experiments, simulations and theory, we provide a quantitative microscopic understanding of the transport and energetics of bacterial rectification—a paradigmatic model of the rectification of living active matter (19, 30). In particular, we identify three essential factors controlling bacterial rectification: (i) a universal distribution of the orientation of bacteria that swim across a linear segment such as the opening of obstacles; (ii) the realignment of bacteria along solid boundaries; (iii) bacterial wobbling. A model incorporating these factors quantitatively describes bacterial transport spanning a wide range of obstacle shapes and predicts the optimal geometry that maximizes the rectification efficiency. Furthermore, we characterize the time irreversibility of bacterial rectification in terms of its local entropy production rate (EPR) and analyze the energetics of the process. Finally, we devise a protocol to measure extractable work through a weak coupling to bacterial flux, which allows us to elucidate the intrinsic relation between three fundamental thermodynamics quantities—particle fluxes, EPR, and extractable work, the first of this kind in a single non-equilibrium process. Taken together, our study provides not only important insights into rectification processes in living systems but also a useful guideline for controlling active particles in biotechnological applications (14, 15, 16, 17, 18).
@@
Away from funnel walls, freely swimming E. coli exhibit the classic run-and-tumble motion (31). The average run length of bacteria in our quasi-2D chamber is significantly longer than the length of the funnel walls. Notably, the trajectory of a bacterium in the run phase is a 3D helix, which manifests as wobbling of the bacterial body under the 2D projection of optical microscopy (32, 33) (Movies 1 and 2). Independent of the angle of incidence, a bacterium always re-aligns and moves parallel to the funnel walls after hitting the walls (Figs. 1A and B) (14). In our minimal simulations, we model bacteria as non-interacting point particles performing run-and-tumble motions without wobbling, and incorporate the surface alignment trait of E. coli via event-driven dynamics (Materials and Methods).
@@
The agreement between our minimal simulations and experiments for $\theta<\theta_{a}$ has two important implications. First, even though bacteria-wall interactions are governed by complex steric and near-field hydrodynamic forces (34, 35, 36, 37), from the perspective of rectification, such complex interactions can be effectively condensed into one simple rule: E. coli reorient themselves parallel to the wall after a collision. Such a simple rule has been ignored in most numerical studies of bacterial rectification (20, 23), with notable exceptions (21, 22). Second, the rectification efficiency can be defined as the net flux per particle, $\eta_{r}\equiv({N}_{+}-{N}_{-})/({N}_{+}+{N}_{-})$. The non-monotonic trend in ${N}_{+}$, in combination with the nearly constant ${N}_{-}$, results in a maximum $\eta_{r}$ at $\theta=\theta_{\max}$ (Fig. 2C).
@@
The hallmark of non-equilibrium processes—with rectification in living systems as a particular example—is the presence of non-zero fluxes and, more fundamentally, time-reversal symmetry breaking (TRSB), measured by the entropy production rate (EPR), $\dot{S}$ (39, 40, 41). EPR gives a direct measure of the minimum energetic cost required to maintain a system out of equilibrium and is crucial in understanding the energetics of non-equilibrium processes.
For a system in non-equilibrium steady state (NESS), the Kullback-Leibler divergence (KLD), $\sigma$, between the probability of observing a time-forward trajectory (X) and its time-reversed counterpart ($\textbf{X}^{\textbf{R}}$) bounds $\dot{S}$ as (42, 43, 44, 45, 46, 47, 48, 49, 23, 43),
@@ Discussion
A recent study has shown that carnivorous plants of genus Genlisea rectify bacterial motion using their root hairs arranged in the shape of funnels. This anatomical structure creates a flux of soil-dwelling bacteria towards the digestive vesicles of the plants (29). Interestingly, the funnel angle of the root hairs fall in the range $90\degree-140\degree$ (29, 54), centering around the optimal rectification angle at $\sim 120\degree$ predicted by our microscopic model and verified by our experiments. This suggests that Genlisea could have structurally evolved to achieve optimal rectification, thereby improving their fitness in nutrient-poor soil conditions.
Our study explores the rectification of bacteria in the dilute limit, where the motions of bacteria are uncorrelated. How do the concepts described here extend to dense systems with correlated bacterial motions? A particularly interesting class of such systems are those displaying active turbulence (55, 56). While it has recently been shown that asymmetric obstacles placed in active turbulent baths can move persistently as autonomous engines (12, 9) and that the statistics of Lagrangian tracers in active turbulent baths carry information about time-irreversibility (57), a unifying thermodynamic framework for active turbulence is still missing, which is an interesting topic for future research.
Is the relation between local fluxes, time irreversibility, and extractable work uncovered in our study applicable to other non-equilibrium systems? Although statistical irreversibility has been widely used to quantify TRSB in various non-equilibrium phenomena (48, 49, 23, 58, 59), simultaneous measurements of time irreversibility, fluxes, and extractable work had not been achieved in any single non-equilibrium systems heretofore. We hope our results will stimulate future research in this direction, which would be of great value for illustrating general principles of non-equilibrium thermodynamics.
@@
A wild-type E. coli strain (BW 25113) genetically modified to express proteorhodopsin (a light-driven transmembrane proton pump) was used in our experiments (60). The bacterial stock was first inoculated in Terrific Broth (TB) supplemented with ampicillin in an orbital shaker at $35\degree$C for 15 hours. This culture was further diluted (1:100) in fresh TB and grown at $30\degree$C for 6-7 hours, with the addition of 1 mM isopropyl $\beta$-d-1-thiogalactopyranoside and 10 $\mu$M methanolic all-trans-retinal in the mid-log phase to trigger the expression of proteorhodopsin. The bacterial suspension was centrifuged ($800g$ for 5 min) to collect motile cells and the supernatant was resuspended in motility buffer (0.01 M potassium phosphate, 0.067 M NaCl, $10^{-4}$ M EDTA, pH 7.0). The suspension was further washed twice, and the final concentration was measured using a biophotometer via optical density at 600 nm (OD600). Bacterial concentrations in all our experiments were fixed below the threshold concentration for collective bacterial swimming (60, 55). With the supply of oxygen, the average swimming speed of bacteria was 25 $\mu$m/s.
@@
The length, width, and gap of the funnel rectifier were fixed at $30$ $\mu$m, $5$ $\mu$m, and $5$ $\mu$m, respectively. The channel depth was kept at 5-6 $\mu$m, so that bacteria were always close to the focal plane set at the middle of the channels. Strong background flows were often observed in the thin channels due to small variations in the channel height inevitably introduced during fabrication. To combat this adverse effect, a water bridge connecting the inlet and the outlet of the channels was established, which equalizes the hydrostatic pressure difference. Polystyrene colloids of 1 $\mu$m in diameter were introduced as test particles in the channels, which showed Brownian diffusion without any detectable drift, ensuring the effectiveness of the procedure. A bacterial suspension of concentration $8\times 10^{7}$ cells/ml was carefully introduced in the microchannels and imaged via an inverted bright-field microscope (Nikon Ti-Eclipse) using a 40×40$\times$40 × objective lens with numerical aperture (NA) 0.65. Videos were acquired at a frame rate of 20 frames per second using an Andor Zyla sCMOS camera. Bacteria were manually counted at the funnel tip for each funnel angle.